Paul Hunt is considering two business ventures. The anticipated returns (in thousands of dollars) of each venture are described by the probability distributions:

Venture A

Earnings

Probability

-20

0.4

40

0.2

50

0.4


Venture B

Earnings

Probability

-15

0.2

30

0.4

40

0.4


Compute (in dollars) the mean and variance for each venture.
Answer #1

.Answer Given venture A Fanningsen probability -20 40 10.4 Mean - Σχε Pi -2000 4) +40 (0-2) + 50(0:4) -8+8+ 20 meam (M = $20variance 0²= {upe prap) - 2 os E1532(0-2) +Bo%C0149 +6803643- 25% - 45 + 360 +670 — 626 = 1045 - 625 To = $420 Variance